### v1

@@ Abstract
Despite the recent advances of the artificial intelligence, building social intelligence remains a challenge. Among social signals, laughter is one of the distinctive expressions that occurs during social interactions between humans. In this work, we tackle a new challenge for machines to understand the rationale behind laughter in video, Video Laugh Reasoning. We introduce this new task to explain why people laugh in a particular video and a dataset for this task. Our proposed dataset, [Uncaptioned image] SMILE, comprises video clips and language descriptions of why people laugh. We propose a baseline by leveraging the reasoning capacity of large language models (LLMs) with textual video representation. Experiments show that our baseline can generate plausible explanations for laughter. We further investigate the scalability of our baseline by probing other video understanding tasks and in-the-wild videos. We release our dataset, code, and model checkpoints on https://github.com/SMILE-data/SMILE.
@@ 2 Related Work
Laughter plays a key role in social interactions, such as bonding, agreement, affection, and emotional regulation Scott et al. (2014). Given its importance in social interactions, seminar works tackle to detect laugh-inducing moments, specifically focusing on humor or sarcasm. Several methods Annamoradnejad and Zoghi (2020); Weller and Seppi (2020) rely primarily on transcripts for humor detection. As laughter occurs with multimodal information, such as variations in tone or facial cues, there are attempts to incorporate audio and text cues from videos Bertero and Fung (2016); Alnajjar et al. (2022), or even include visual cues Castro et al. (2019); Hasan et al. (2019); Ray et al. (2022) to pinpoint the occurrences of humor. Yet they focus on detecting whether a certain situation induces laughter or predicting the intensity of laughter, without providing explanations for the underlying reasons behind the laughter (See Figure 1). Moreover, despite the availability of datasets for understanding the types and characteristics of laughing moments Urbain et al. (2010); McKeown et al. (2012); Dupont et al. (2016), no dedicated dataset is available for comprehending the context surrounding laughter. Few works Chowdhery et al. (2022); Hessel et al. (2023) have attempted to reason about laughter or jokes. However, their scope differs from ours, as they focus on providing instant textual descriptions of humor or cartoon images accompanied by text. To the best of our knowledge, we are the first to introduce the task of understanding the reason for laughter within videos, accompanied by our comprehensive dataset.
@@ 3.3 Data Analysis
4 Experiment
[…]
@@ 4 Experiment
[…]
@@ A B A wins (%) Fleiss’-$\kappa$
We split our dataset into 5 cross-validation splits except for the test set. We use the training set either for fine-tuning or as a few-shot context for the language model and test set for evaluation. Implementation details are in Appendix D.
Evaluation metrics
[…]
4.1 Comparison with video model
@@ 4.2 Evaluation
Figure 4: Qualitative results on laugh reasoning. For the examples in (a), GPT3 Brown et al. (2020a) fine-tuned on our dataset (denoted FT w/ A+V+T) understands the reasons for laughter by referencing multimodal cues. In contrast, the model fine-tuned using the transcript-only (denoted FT w/ T) manages to understand the reasons partially. The visual cues (scene description) are crucial for capturing “joey’s sudden appearance” which is important to infer the reason for laughter in (b).
We analyze our baseline on the laugh reasoning task in various setups. We utilize both quantitative metrics and human evaluation. Quantitative results are in Table 3, and the results and human agreement are in Table 4. Our evaluations aim to address four key questions:
@@
Q4. Does the model explain the reason for laughter as well as humans?   No, the human-annotated laughter reasons are preferred by 66.2% than those generated by fine-tuned GPT-3 (our best model) as shown in Q4 of Table 4. Figure 4 (b) provides an example illustrating the comparison between human-annotated reasons (GT) and generated reason for laughter. In this sample, all crowd workers prefer GT because the model struggles to distinguish the subtle difference between surprise and posed surprise.
@@ 5.1 Evaluation on other tasks
We follow the same train/test split, and evaluation procedure as in the benchmark measuring accuracy as the percentage of predicted values matching with GT. We use LLaMA and GPT-3 for training with textual representation converted from the video in the training set of each benchmark dataset. Table 5 shows that our method achieves strong performance [^4] We do not compare with FunnyNet Liu et al. (2022) as they use an additional large-scale dataset for training. on both tasks. This experiment highlights the scalability of utilizing LLMs with textual representation in various video understanding tasks.
@@ 5.2 Evaluation on the in-the-wild videos
We extend our laughter reasoning to in-the-wild videos, encompassing different video types and laugh contexts compared to our dataset. First, we evaluate our approach on a video clip from a standup comedy, which has similar audience laughter patterns to those in our dataset. We convert the video into textual representation and infer the reason for the audience laughing. Figure 5 shows that the model can plausibly generate an explanation for the reason for laughter in standup comedy.
Next, we test on a video clip featuring an intimate conversation between a married couple. In this case, the laughter originates from the speakers themselves rather than the audience. As this does not belong to the comedic genre but rather a sincere conversation between two people, it is more likely that non-humor-based laughter, such as nervous or social laughter, may occur. Figure 5 shows that the model can also understand the nervous laughter used to alleviate tension or awkwardness in the situation.
@@ 6 Conclusion
In this paper, we aim to understand the reason behind laughter by introducing Video Laugh Reasoning task, accompanied with [Uncaptioned image] SMILE dataset. While the model did not surpass human capabilities, we show that the model can generate plausible explanations about laughter reason, underlining that multimodal cues in our dataset nicely infuse the laugh reasoning capacity to the model. We also show the results applied to other tasks and other types of video, hinting at the scalability of utilizing LLM with multimodal textual representation.
@@
Our work can be regarded as a stepping stone toward developing socially intelligent agents that understand and appropriately create non-verbal cues, such as laughter, playing a crucial role in building rapport, expressing emotions, and creating deep emotional exchanges Tickle-Degnen and Rosenthal (1990); Argyle (1972). This enables expressive interactions with humans beyond the existing dialogue agents, e.g., ChatGPT and Amazon Alexa, which mostly focus on verbal signals.
Furthermore, our work shows the possibility of formulating abstract concepts in human interaction, such as the reason for laughter, by leveraging the capacity of LLMs with textual representation. It is promising to apply our method to solve other complex problems, such as deciphering nuanced body language in diverse cultures, predicting audience reactions or engagement levels to media contents or advertisements, and detecting subtle signs of cognitive or emotional disorders. These applications could provide insights into intercultural communication, marketing, and health diagnostics.
@@ Appendix D Implementation details
GPT3 zero-shot and in-context learning
We use the official GPT3 Brown et al. (2020a), a non-free commercial version, as follows. We utilize the davinci-text-002 model of GPT3 Brown et al. (2020a) for the zero-shot and in-context learning experiments. Examples of the prompts for both tasks are shown in Figure 11. The “prompt” provides the context of the task and the multimodal cues of the video, and “completion” provides the reason for the laughter. The zero-shot setup only takes “prompt” and generates the reason for the laughter, while the in-context learning setup is given with additional three randomly labeled samples from the training set as few-shot examples.
@@
[…]
@@
[…]

### v2

@@ Abstract
Despite the recent advances of the artificial intelligence, building social intelligence remains a challenge. Among social signals, laughter is one of the distinctive expressions that occurs during social interactions between humans. In this work, we tackle a new challenge for machines to understand the rationale behind laughter in video, Video Laugh Reasoning. We introduce this new task to explain why people laugh in a particular video and a dataset for this task. Our proposed dataset, [Uncaptioned image] SMILE, comprises video clips and language descriptions of why people laugh. We propose a baseline by leveraging the reasoning capacity of large language models (LLMs) with textual video representation. Experiments show that our baseline can generate plausible explanations for laughter. We further investigate the scalability of our baseline by probing other video understanding tasks and in-the-wild videos. We release our dataset, code, and model checkpoints on https://github.com/postech-ami/SMILE-Dataset.
@@ 2 Related Work
Laughter plays a key role in social interactions, such as bonding, agreement, affection, and emotional regulation Scott et al. (2014). Given its importance in social interactions, seminar works tackle to detect laugh-inducing moments, specifically focusing on humor or sarcasm. Several methods Annamoradnejad and Zoghi (2020); Weller and Seppi (2020) rely primarily on transcripts for humor detection. As laughter occurs with multimodal information, such as variations in tone or facial cues, there are attempts to incorporate audio and text cues from videos Bertero and Fung (2016); Alnajjar et al. (2022), or even include visual cues Castro et al. (2019); Hasan et al. (2019); Ray et al. (2022) to pinpoint the occurrences of humor. Yet they focus on detecting whether a certain situation induces laughter or predicting the intensity of laughter, without providing explanations for the underlying reasons behind the laughter (See Figure 1). Moreover, despite the availability of datasets for understanding the types and characteristics of laughing moments Urbain et al. (2010); McKeown et al. (2012); Dupont et al. (2016), no dedicated dataset is available for comprehending the context surrounding laughter. Few works Chowdhery et al. (2022); Hessel et al. (2023); Ko et al. (2023) have attempted to reason about laughter or jokes. However, their scope differs from ours, as they focus on providing instant textual descriptions of humor or cartoon images accompanied by text. To the best of our knowledge, we are the first to introduce the task of understanding the reason for laughter within videos, accompanied by our comprehensive dataset.
@@ 3.3 Data Analysis
Model BLEU4 ($\uparrow$) METEOR ($\uparrow$) ROUGEL ($\uparrow$) BERTScore ($\uparrow$) Win rate
@@
4 Experiment
Model Num. of parameters Modality BLEU4 ($\uparrow$) METEOR ($\uparrow$) ROUGEL ($\uparrow$) BERTScore (F1) ($\uparrow$)
@@ 4 Experiment
We split our dataset into 5 cross-validation splits except for the test set. We fine-tune two LLMs, GPT-3 Brown et al. (2020a) and LLaMA Touvron et al. (2023) with the training set and use the test set for evaluation.
Implementation details
We use the official GPT-3 Brown et al. (2020a), a non-free commercial version, as follows. We utilize the davinci-text-002 model of GPT-3 Brown et al. (2020a) for the zero-shot and in-context learning experiments. Examples of the prompts for both tasks are shown in Figure 4. The “prompt” provides the context of the task and the multimodal cues of the video, and “completion” provides the reason for the laughter. The zero-shot setup only takes “prompt” and generates the reason for the laughter, while the in-context learning setup is given with additional three randomly labeled samples from the training set as few-shot examples. More implementation details including LLaMA are in Appendix D.
Evaluation metrics
We utilize both quantitative metrics and human evaluation. We use metrics commonly employed for evaluating language generation tasks, including BLEU4 Papineni et al. (2002), METEOR Banerjee and Lavie (2005), ROUGEL Lin (2004), and BERTscore Zhang et al. (2019). For the human evaluation, we gather assessments from 3 crowd-workers per test sample by asking them to select their preferred explanation for laughter from a pair of options and take a majority vote to determine a winner. We calculate the average win rate (%) over the test set.
4.1 Comparison with video model
@@ 4.2 Evaluation
Figure 5: Qualitative results on laugh reasoning. For the examples in (a), GPT-3 Brown et al. (2020a) fine-tuned on our dataset (denoted FT w/ A+V+T) understands the reasons for laughter by referencing multimodal cues. In contrast, the model fine-tuned using the transcript-only (denoted FT w/ T) manages to understand the reasons partially. The visual cues (scene description) are crucial for capturing “joey’s sudden appearance” which is important to infer the reason for laughter in (b).
We analyze our baseline on laugh reasoning in various setups. We utilize both quantitative and human evaluation. Quantitative results are in Table 3, and the results of human agreements are in Table 4. Our evaluations aim to address four key questions.
@@
Q4. Does the model explain the reason for laughter as well as humans?   No, the human-annotated laughter reasons are preferred by 66.2% than those generated by fine-tuned GPT-3 (our best model) as shown in Q4 of Table 4. Figure 5 (b) provides an example illustrating the comparison between human-annotated reasons (GT) and generated reason for laughter. In this sample, all crowd workers prefer GT because the model struggles to distinguish the subtle difference between surprise and posed surprise, while the human-annotated reason successfully captures it.
@@ 5.1 Evaluation on other tasks
We follow the same train/test split, and evaluation procedure as in the benchmark for measuring the accuracy of each detection task. We use LLaMA and GPT-3 for training with textual representation converted from the video in the training set of each benchmark dataset. Table 5 shows that our method achieves strong performance [^4] We do not compare with FunnyNet Liu et al. (2022) as they use an additional large-scale dataset for training. on both tasks. This experiment highlights the scalability of utilizing LLMs with textual representation in various video understanding tasks.
@@ 5.2 Evaluation on the in-the-wild videos
We extend our laughter reasoning to in-the-wild videos, encompassing different video types and laughter contexts compared to our dataset. First, we evaluate our approach on a video clip from a stand-up comedy, which has similar audience laughter patterns to those in our dataset. We convert the video into a textual representation and infer the reason for the audience laughing. Figure 6 shows that the model can generate a plausible explanation for the reason for laughter in stand-up comedy.
Next, we test on a video clip featuring an intimate conversation between a married couple. In this case, the laughter originates from the speakers themselves rather than from the audience. As this does not belong to the comedic genre but rather a sincere conversation between two people, it is more likely that non humor-based laughter, such as nervous or social laughter, may occur. Figure 6 shows that the model can also understand the nervous laughter used to alleviate tension or awkwardness in the situation.
@@ 6 Conclusion
In this paper, we aim to understand the reason behind laughter by introducing Laugh Reasoning task, accompanied with SMILE dataset. While the model did not surpass human capabilities, we show that the model can generate plausible explanations about laughter reason, underlining that multimodal cues in our dataset nicely infuse the laugh reasoning capacity to the model. We also show the results applied to other tasks and other types of video, hinting at the scalability of utilizing LLM with multimodal textual representation.
@@
Our work can be regarded as a stepping stone toward developing socially intelligent agents that understand and appropriately create non-verbal cues, such as laughter, playing a crucial role in building rapport, expressing emotions, and creating deep emotional exchanges Tickle-Degnen and Rosenthal (1990); Argyle (1972). Such advancement moves us beyond the capabilities of current dialogue agents, e.g., ChatGPT or Alexa, which mostly focus on verbal signals. Incorporating 3D talking head methods Sung-Bin et al. (2024); Zhao et al. (2024) could offer the way agents are visualized, enabling more expressive and multimodal interactions with users.
Acknowledgement
This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (No.2022-0-00290, Visual Intelligence for Space-Time Understanding and Generation based on Multi-layered Visual Common Sense and No.2022-0-00124, Development of Artificial Intelligence Technology for Self-Improving Competency-Aware Learning Capabilities and No.2021-0-02068, Artificial Intelligence Innovation Hub) and NCSOFT.
@@ Appendix D Implementation details
Test dataset Train dataset Modality BLEU4 ($\uparrow$) METEOR ($\uparrow$) ROUGEL ($\uparrow$) BERTScore (F1) ($\uparrow$)
@@
Test dataset Train dataset Modality BLEU4 ($\uparrow$) METEOR ($\uparrow$) ROUGEL ($\uparrow$) BERTScore (F1) ($\uparrow$)